• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Yi Qiang

    Log entry time 16:04:31 on March 15, 2009

    Entry number 266325

    keyword=Day Shift Summary

    Day Shift Summary:

    Shift Workers
    -----------------------------------------------------------------------------------
    SL: Yi Qiang
    TO: Kalyan Allada

    Shift Summary
    -----------------------------------------------------------------------------------
    08:00 Shift starts with production of Kin 4H
    08:19 BB Run 2181 stopped due to ROC10 disconnection
    08:38 BB DAQ finally became normal after 3 reboots
    10:00 Switch to Kin 4F
    10:06 NMR: 60.65%
    11:05 Target to N2 reference cell: 116 psig
    11:39 Target back to 3He target for production
    14:14 IOC IOCHLA2 for Hall A raster is down
    14:43 IOC fixed, no BPM readings
    14:46 All problems fixed, continue production with raster size 4x5.2 MCC
    value
    15:35 Increase raster size to 6x5.2 MCC value
    15:54 BigBite HV tripped
    16:00 Shift ends with BigBite HV recovering

    Run Summary
    -----------------------------------------------------------------------------------
    BigBite
    07:27 - 08:19 #2181 3He 15uA Prod. 6.3M Stopped due to 100% dead time,
    End of run failed
    08:24 - 08:26 #2182 3He 15uA Junk.
    08:38 - 09:30 #2183 3He 15uA Prod. 6.5M
    09:32 - 10:04 #2184 3He 15uA Prod. 4.12M
    10:08 - 10:59 #2185 3He 15uA Prod. 6.5M
    11:05 - 11:37 #2186 N2 5uA Calib. 4.13M
    11:43 - 12:32 #2187 3He 15uA Prod. 6.5M
    12:33 - 13:14 #2188 3He 15uA Prod. 4.83M
    13:18 - 14:09 #2189 3He 15uA Prod. 5.43M Stopped due to MCC IOC problem
    14:46 - 15:34 #2190 3He 15uA Prod. 5.3M
    15:37 - 15:54 #2191 3He 15uA Prod. 2.1M Bigbite HV tripped

    HRSL
    Kin 4H:
    07:27 - 09:30 #20699 3He 15uA Prod. 534k 0.1112C
    09:32 - 09:55 #20700 3He 15uA Prod. 83k 0.0176C
    Kin 4F
    10:09 - 11:00 #20701 3He 15uA Prod. 279k 0.0458C
    11:05 - 11:37 #20702 N2 5uA Calib. 208k *0.00939C*
    11:43 - 13:14 #20703 3He 15uA Prod. 487k 0.0801C
    13:18 - 14:10 #20704 3He 15uA Prod. 247k 0.0379C Stopped due to MCC IOC
    problem
    14:47 - 15:34 #20705 3He 15uA Prod. 229k 0.0365C
    15:37 - #20706 3He 15uA Prod.

    Compton/Happex
    07:39 - 09:39 #16433/31035/60813
    09:41 - 11:42 #16434/31036/60814
    12:08 - 12:55 #16435/31037/60815
    12:55 - 13:26 #16436/31037/60815
    13:27 - 14:10 #16436/31038/60816
    14:51 - #16437/31039/60817